MULTI-WATT, SINGLE FREQUENCY OUTPUT FROM 1.4 TO 4.2 MICRONS FROM A SINGLE LASER SYSTEM

A. HENDERSON, R. STAFFORD, and P. HOFFMAN, Aculight Corporation, 22121 20th Ave SE Bothell, WA 98021-4408, USA..

The fiber-laser-pumped continuous-wave optical parametric oscillator^{a}. (CW OPO) is an ultra-versatile laser system for high resolution near to midinfrared spectroscopy. These devices are currently being offered as a commercial product by Aculight Corporation. We will describe the operating characteristics of the OPO system which provides up to 5 Watts single frequency output between 1.4 and 2 microns and up to 4 Watts between 2.3 and 4.2 microns. Key characteristics for spectroscopists are a mode-hop free tuning range of 60GHz with kHz scan rates and a measured linewidth of 30kHz. In addition we will provide details of long term amplitude and frequency stability, beam quality and further lower power single frequency outputs in the visible range.

aA. Henderson and R. Stafford, Opt. Express 14, 767 (2006).